OpenStudy (anonymous):

need to find t such that (t, -7/9) is a point on the unit circle.

OpenStudy (anonymous):

you'll need to solve this equation: t^2 + (-7/9)^2 = 1. since this is a quadratic in terms of t, you'll get two answers... lemme know if you need help.

OpenStudy (anonymous):

i can get the answer, but what is -7/9 squared?

OpenStudy (anonymous):

-7/9 * -7/9 = 49/81

OpenStudy (anonymous):

ty haha

OpenStudy (anonymous):

yw